# Designed self-assembly of programmable colloidal atom-electron equivalents

### v2

@@ Abstract
We aim to guide the targeted binding between $\alpha$ and $\gamma$ type PAEs through EE-mediated interactions. However, as shown in Fig. 1(c), due to the multivalency of EEs, non-targeted interactions exist between particles of the same type. Our goal is to understand how to design these interactions to optimize the ability to form targeted structures. We first perform grand canonical Monte Carlo (GCMC) simulations incorporating explicit EEs to validate the pair interactions. In our simulations, PAEs are modeled as mobile DNA-coated colloids (mDNACCs), while the EEs are treated as zero-volume mDNACCs that can interact with both $\alpha$ and $\gamma$ types. The interactions within the system are calculated based on the framework of reversible mobile binders proposed in Ref. Angioletti-Uberti et al. (2014). Two PAEs of the same ($\alpha\alpha$) or different types ($\alpha\gamma$) are placed at fixed positions within a simulation box, while the EEs are inserted and deleted using standard GCMC moves Frenkel and Smit (2023). The effective interactions are calculated for a given chemical potential of EEs $\mu$ and a specified center-to-center distance $r$ between two PAEs using the thermodynamic integration (see SM Sec. III sup ). As shown in Fig. 2 (a,b), we plot the pair interactions between the same and different types of PAEs, $\beta U_{\rm eff}^{\alpha\alpha}$ and $\beta U_{\rm eff}^{\alpha\gamma}$, demonstrating that our method quantitatively agrees with the explicit EE simulations without requiring any fitting parameter. Our method captures the number of unbound ssDNAs on PAEs, i.e., $\bar{n}^{\alpha\gamma}/n$ with $n_{\alpha}=n_{\gamma}=n=100$, and the number of EEs between two PAEs, i.e., $N_{E}^{\alpha\gamma}$, and these results also agree quantitatively with the simulation data (Fig. 2(c,d)). Furthermore, our method provides detailed information on EEs in different substates.
Entropy driven effective PAE-PAE interactions.
@@
Similar to the bivalent linker-mediated DNACCs and polymer-microemulsion networks Xia et al. (2020); Zilman et al. (2003), as shown in Fig. 3(a,b), we observe that the strength of effective colloidal interactions $\beta U_{\rm eff}^{\alpha\alpha/\alpha\gamma}(r^{*}=2R)$ reaches a plateau at the strong binding energy or low temperature limit ($\beta\Delta G\to-\infty$). The reason is that at this limit, $\bar{n}_{i}\to 0$, and all ssDNAs on any PAE $i$ are occupied. Therefore, for any PAE configurations, the energy part remains the same as the configuration where all PAEs are isolated, and has no contribution to the effective interaction $\beta U_{\rm eff}$. At this limit, $\beta U_{\rm eff}$ solely depends on the combinatorial and configurational entropies, i.e., the valency and chemical potential of EEs. Based on the additional constraint $\bar{n}_{i}\to 0$ and taking the reference state as $\mathcal{A}^{1}$, we introduce the entropy driven effective binding strength of an EEs under $\mathcal{A}_{i}^{k}$ substate, $\tilde{\Xi}_{\mathcal{A},i}^{k}$, and under $\mathcal{B}_{ij}^{kl}$ substate, $\tilde{\Xi}_{\mathcal{B},ij}^{kl}$, respectively (see SM Sec. II sup )
$\tilde{\Xi}_{\mathcal{A},i}^{k}=$ $\left.{\Xi_{\mathcal{A},i}^{k}}\middle/{\left(\Xi_{\mathcal{A},i}^{1}\right)^{k}}\right.=\frac{\binom{\lambda}{k}m_{i}^{0}}{\left(m_{i}^{0}\lambda\right)^{k}},$ (6a)
$\tilde{\Xi}_{\mathcal{B},ij}^{kl}=$ $\left.{\Xi_{\mathcal{B},ij}^{kl}}\middle/\left[{\left(\Xi_{\mathcal{A},i}^{1}\right)^{k}\left(\Xi_{\mathcal{A},j}^{1}\right)^{l}}\right]\right.=\frac{w^{kl}m_{ij}^{0}}{\left(m_{i}^{0}\lambda\right)^{k}\left(m_{j}^{0}\lambda\right)^{l}},$ (6b)
@@
To demonstrate the entropy driven PAE-PAE interactions in many-body systems, we perform $NVT$ MC simulations with the effective interaction from Eq. 1 and 5 for an equimolar binary mixture of PAEs at the packing fraction $\eta=4N\pi R^{3}/3V=0.094$ with various $\beta\Delta G$ and $\lambda$, where we consider a PAE as a singlet if there is no bridge formed with other PAEs (see SM Sec. III sup ). As shown in Fig. 3 (c), the singlet fraction decreases and reaches a plateau with decreasing $\beta\Delta G$, following the same trend as the pair interaction (Fig 3(a,b)).
Anomalous EE valency dependence.
@@
As shown in Fig. 3(a,b), approaching the strong binding limit ($\beta\Delta G\to-\infty$), intriguingly, the effective interaction strength between the same type of PAEs, $\beta U_{\rm eff}^{\alpha\alpha}(r^{*})$, becomes stronger as the valency of EE $\lambda$ increases, while the effective interaction between different types of PAEs, $\beta U_{\rm eff}^{\alpha\gamma}(r^{*})$, shows the opposite dependence. To clarify these trends, we plot $\beta U_{\rm eff}^{\alpha\alpha/\alpha\gamma}(r^{*})$ as functions of $\lambda$ at the strong binding limit, shown as the black solid and dashed curves in Fig. 4. We note that at this limit, $\bar{n}$ is almost independent of distance, i.e., the number of bridging EEs $Q$ is small enough, and the effective pair interaction $\beta U_{\rm eff}$ mainly depends on $Q$, with $\beta U_{\rm eff}\approx-Q$. For the $\alpha\gamma$ interaction, we have the relation ${Q_{ij}}/{m_{ij}^{0}}={M_{i}M_{j}}/{m_{i}^{0}m_{j}^{0}}$, i.e., $Q_{ij}\sim M_{i}M_{j}$, indicating that bridging is a joint event depending on the adsorption on both particles, where adsorption promotes bridging. Therefore, as $\lambda$ increases, both the numbers of EEs in the adsorbed state and bridging state decrease, leading to a weaker effective EE-mediated interaction. Differently, for the $\alpha\alpha$ interaction, we have $Q_{ij}\sim-M_{i}-M_{j}$, indicating that bridging competes with adsorption. Thus, increasing the $\lambda$ decreases the number of EEs in the adsorbed state while increasing the number of EEs in the bridging state, resulting in a stronger effective PAE interaction.
@@
Furthermore, we observe that with increasing $\lambda$, indicated by the gray horizontal line in Fig. 4, the effective PAE interactions unexpectedly converge and become independent of the binding energy and the type of pairs. At this limit, we also have $\bar{n}_{i}\to 0$, similar to the strong binding limit according to Eq. 4. Thus, $\beta F_{\rm att}$ in Eq. 7 holds with the effective binding strength (see SM Sec. II sup )
@@
This indicates that increasing the valency has less and less impact on effective PAE interactions.
Figure 4: Anomalous $\lambda$ dependence of effective interaction. The strength of EE-mediated interactions between PAEs of the same type ($\beta U_{\rm eff}^{\alpha\alpha}$, dashed curves and open symbols) and the different types same type($\beta U_{\rm eff}^{\alpha\gamma}$, solid curves and solid symbols) as functions of EE valency $\lambda$ with different binding energy $\beta\Delta G$. Symbols are calculated from numerical simulations. The vertical dotted lines show the location of optimal valency $\lambda^{*}$ where $|\beta U_{\rm eff}^{\alpha\gamma}(r^{*})-\beta U_{\rm eff}^{\alpha\alpha}(r^{*})|$ maximizes at different $\beta\Delta G$. The horizontal grey line indicates $\beta U_{\rm eff}(r^{*})$ at large $\lambda$ limit. Parameters: $\beta\mu=-13$, $n=100$, $r^{\ast}=2R$ and $l_{r}=0.1R$.
At finite binding energies, the effect of increasing $\lambda$ depends on the types of PAEs involved. As shown in Fig. 4, for interactions between the same type of PAEs ($\beta U_{\rm eff}^{\alpha\alpha}$), increasing the EE valency leads to stronger interactions that eventually converge, similar to the behavior in the strong binding limit. Differently, the situation is more complex for interactions between different types of PAEs ($\beta U_{\rm eff}^{\alpha\gamma}$). At weak binding energies, e.g., $\beta\Delta G=-10$, increasing $\lambda$ strengthens the PAE-PAE interaction; at strong binding energies, e.g., $\beta\Delta G=-14$, increasing the EE valency weakens the PAE-PAE interaction; and at intermediate binding energies, e.g., $\beta\Delta G=-12$, the dependence of the strength of PAE-PAE interaction on $\lambda$ is non-monotonic. Stronger binding energies cause the effective interaction to more closely resemble that of the strong binding limit, whereas weaker binding energies result in PAE-PAE interactions that are naturally weaker than this limit.
@@
As all interactions converge at large $\lambda$, one can find that there exists an optimal valency $\lambda^{*}$ at which the difference between strengths of effective interactions among different PAEs $|\beta U_{\rm eff}^{\alpha\gamma}(r^{*})-\beta U_{\rm eff}^{\alpha\alpha}(r^{*})|$ maximizes, and $\lambda^{*}$ decreases with decreasing the binding free energy $\Delta G$ (Fig. 4). This optimal $\lambda$ maximizes the disparity between targeted ($\alpha\gamma$) and non-targeted ($\alpha\alpha$) PAE interactions, and our results suggest that at the strong binding limit ($\Delta G\to-\infty$), the optimal valency of EE for mediating specific binding between targeted PAEs is $\lambda^{*}=1$. However, we note that experimentally the binding strength is normally not very strong, as strong binding would make the system hard to reach the designed state in equilibrium.
Discussion.
@@
To conclude, we have studied the self-assembly of a minimal binary system of PAE-EEs, where EEs act as multivalent linkers with two types of ssDNA ends complementary to those on PAEs. We derived an explicit formula for the effective PAE-PAE interaction mediated by EEs with arbitrary valency $\lambda$, which quantitatively agrees with simulations incorporating explicit EEs. Our analysis reveals an anomalous dependence of the effective PAE-PAE interactions on $\lambda$. Specifically, at the strong binding limit, we found that increasing $\lambda$ strengthens the interaction between PAEs of the same type but weakens that between different types, which converges to the same value in the large valency limit. We identified an optimal valency $\lambda^{*}$ that maximizes the effective interaction difference between the targeted and non-targeted PAE pairs, which provides a practical design principle for the designed self-assembly of PAE-EE systems. Moreover, our explicit formula can be directly used in MC simulation to efficiently sample the thermodynamic and structural information of the system. Beyond motivating experimental work towards the design of complex DNA structures, we provide the tools for modeling and analyzing soft matter and biological systems with multivalent linkers, e.g., polyelectrolyte Muthukumar (2017); Yu et al. (2018), hydrogels Ooi et al. (2020); Le Roy et al. (2024), vitrimersRöttger et al. (2017); Lei et al. (2020), multivalent drug design Mammen et al. (1998); Dubacheva et al. (2023), and cell adhesion and signaling Zhu et al. (2008).
Acknowledgements.
@@
X. X. acknowledges support from the Alexander von Humboldt-Stiftung. This work was financially supported by the Academic Research Fund from the Singapore Ministry of Education (RG151/23 and MOE2019-T2-2-010) and the National Research Foundation, Singapore, under its 29th Competitive Research Program (CRP) Call (NRF-CRP29-2022-0002).

### v3

@@ 3.1 Numerical verification
We aim to guide the targeted binding between $\alpha$ and $\gamma$ type PAEs through EE-mediated interactions. However, as shown in Fig. 1(c), due to the multivalency of EEs, EE-mediated interactions exist between particles of the same type and different types. We consider $\beta U_{\rm eff}^{\alpha\gamma/\alpha\alpha}$ to be the targeted/non-targeted PAE-PAE interaction, respectively. Our goal is to understand how to design these interactions to optimize the ability to form targeted structures. We first perform grand canonical Monte Carlo (GCMC) simulations incorporating explicit EEs to validate the pair interactions. In our simulations, PAEs are modeled as mobile DNA-coated colloids (mDNACCs), while the EEs are treated as zero-volume mDNACCs that can interact with both $\alpha$ and $\gamma$ PAEs. The interactions within the system are calculated based on the framework of reversible mobile binders proposed in Ref. [28]. Two PAEs of the same ($\alpha\alpha$) or different types ($\alpha\gamma$) are placed at fixed positions within a simulation box, while the EEs are inserted and deleted using standard GCMC moves [33]. The effective interactions are calculated for a given chemical potential of EEs $\mu$ and a specified center-to-center distance $r$ between two PAEs using the thermodynamic integration (see Supplementary Material Sec. IV.A [32]). As shown in Fig. 2 (a,b), we plot the pair interactions between the same and different types of PAEs, $\beta U_{\rm eff}^{\alpha\alpha}$ and $\beta U_{\rm eff}^{\alpha\gamma}$, demonstrating that our method quantitatively agrees with the explicit EE simulations without any fitting parameter. Our method captures the number of unbound ssDNAs on PAEs, i.e., $\bar{n}^{\alpha\gamma}/n$ with $n_{\alpha}=n_{\gamma}=n=100$, and the number of EEs between two PAEs, i.e., $N_{E}^{\alpha\gamma}$, and these results also agree quantitatively with the simulation results (Fig. 2(c,d)). Furthermore, our method provides detailed information on EEs in different substates.
Figure 2: Numerical verification of the effective pair interactions between PAEs. (a,b) EE-mediated effective PAE-PAE interactions $\beta U_{\rm eff}(r)$ as functions of the center-to-center distance $r$ between two PAEs for different EE valency $\lambda$, where the PAEs are of the same type (a, $\beta U_{\rm eff}^{\alpha\alpha}$) and different types (b, $\beta U_{\rm eff}^{\alpha\gamma}$). Note a negative effective potential $\beta U_{\mathrm{eff}}$ corresponds to an attractive interaction. (c) The ratio of unbound ssDNAs $\bar{n}^{\alpha\gamma}/n$ as a function of $r$ for different $\lambda$ between PAEs of different types. (d) Number of EEs $N_{\rm E}^{\alpha\gamma}$ depicted by the shadow area in the inset as a function of $r$ for different $\lambda$ between PAEs of different types. Symbols are from numerical simulations, and curves are theoretical predictions. Parameters: $\beta\mu=-14$, $\beta\Delta G=-10$, $n=100$ and $l_{r}=0.1R$.
Figure 3: Entropy driven PAE-PAE interactions. (a,b) The strength of EE-mediated interaction $\beta U_{\rm eff}(r^{*})$ as functions of individual binding energy $\beta\Delta G$ between two PAEs for different EE valency $\lambda$ at $\beta\mu=-13$, where the PAEs are of the same type (a, $\beta U_{\rm eff}^{\alpha\alpha}(r^{*})$) and different types (b, $\beta U_{\rm eff}^{\alpha\gamma}(r^{*})$). Symbols are from numerical simulations, and curves are theoretical predictions. (c) percentage of unbridged PAEs in manybody PAE-EEs systems at $\eta=0.094$ as a function of $\beta\Delta G$ for different $\lambda$ at $\beta\mu=-11$. Symbols are obtained by the many-body Monte Carlo simulations, and the dashed lines are guided by eyes. Parameters: $n_{\alpha}=n_{\gamma}=100$, $r^{\ast}=2R$ and $l_{r}=0.1R$.
3.2 Entropy driven effective PAE-PAE interactions
@@
As shown in Fig. 3(a,b), similar to the bivalent linker-mediated DNACCs and polymer-microemulsion networks [16, 34], we observe that the strength of effective colloidal interactions at contact $\beta U_{\rm eff}^{\alpha\alpha/\alpha\gamma}(r^{\ast})$ saturates and does not diverge at the strong binding or low temperature limit under any valency of EEs. The reason is that at the strong binding limit, all ssDNAs on any PAE $i$ are occupied, i.e., $\bar{n}_{i}\to 0$, therefore, the energy remains the same as the configuration where all PAEs are isolated for any PAE configurations. As shown in Eq. 1, the energy term cancels out in $\beta F_{\rm att}$ and $\beta F_{\rm att}^{\prime}$ and therefore does not contribute to the effective interaction $\beta U_{\rm eff}$. At this limit, $\beta U_{\rm eff}$ solely depends on the combinatorial and configurational entropies, i.e., the valency and chemical potential of EEs. Based on the additional constraint $\bar{n}_{i}\to 0$ and taking $\mathcal{A}^{1}$ as the reference state, we introduce the entropy driven effective binding strength of an EEs under $\mathcal{A}_{i}^{k}$ substate, $\tilde{\Xi}_{\mathcal{A},i}^{k}$, and under $\mathcal{B}_{ij}^{kl}$ substate, $\tilde{\Xi}_{\mathcal{B},ij}^{kl}$, respectively (see Supplementary Material Sec. III.A [32])
$\tilde{\Xi}_{\mathcal{A},i}^{k}=$ $\left.\frac{\Xi_{\mathcal{A},i}^{k}}{\left(\Xi_{\mathcal{A},i}^{1}\right)^{k}}\right.=\frac{\binom{\lambda}{k}m_{i}^{0}}{\left(m_{i}^{0}\lambda\right)^{k}},$ (11a)
$\tilde{\Xi}_{\mathcal{B},ij}^{kl}=$ $\left.\frac{\Xi_{\mathcal{B},ij}^{kl}}{\left[{\left(\Xi_{\mathcal{A},i}^{1}\right)^{k}\left(\Xi_{\mathcal{A},j}^{1}\right)^{l}}\right]}\right.=\frac{w^{kl}m_{ij}^{0}}{\left(m_{i}^{0}\lambda\right)^{k}\left(m_{j}^{0}\lambda\right)^{l}},$ (11b)
@@
To demonstrate the entropy driven PAE-PAE interactions in many-body systems, we perform $NVT$ MC simulations with the effective interaction from Eq. 1 and 9 for an equimolar binary mixture of PAEs at the packing fraction $\eta=4N\pi R^{3}/3V=0.094$ with various $\beta\Delta G$ and $\lambda$ (see Supplementary Material Sec. IV.B [32]). As shown in Fig. 3 (c), the percentage of unbridged PAEs decreases and reaches a plateau with decreasing $\beta\Delta G$, following the same trend as the pair interaction (Fig 3(a,b)). Additional many-body simulations at higher packing fractions ($\eta=$ 0.094–0.282, Fig. S1) show the same unbridged-fraction trends, confirming that our conclusions are insensitive to the PAE concentration.
Figure 4: Anomalous $\lambda$ dependence of effective interaction. The strength of EE-mediated interactions between PAEs of the same type ($\beta U_{\rm eff}^{\alpha\alpha}$, dashed curves and open symbols) and the different types ($\beta U_{\rm eff}^{\alpha\gamma}$, solid curves and solid symbols) as functions of EE valency $\lambda$ with different binding energy $\beta\Delta G$. Symbols are calculated from numerical simulations. The vertical dotted lines show the location of optimal valency $\lambda^{*}$ where $|\beta U_{\rm eff}^{\alpha\gamma}(r^{*})-\beta U_{\rm eff}^{\alpha\alpha}(r^{*})|$ maximizes at different $\beta\Delta G$. The horizontal grey line indicates $\beta U_{\rm eff}(r^{*})$ at large $\lambda$ limit. Parameters: $\beta\mu=-13$, $n=100$, $r^{\ast}=2R$ and $l_{r}=0.1R$.
3.3 Anomalous EE valency dependence
@@
As shown in Fig. 3(a,b), approaching the strong binding limit ($\beta\Delta G\to-\infty$), intriguingly, the effective interaction strength at contact between the same type of PAEs, $\beta U_{\rm eff}^{\alpha\alpha}(r^{*})$, becomes stronger as the valency of EE $\lambda$ increases, while the effective interaction at contact between different types of PAEs, $\beta U_{\rm eff}^{\alpha\gamma}(r^{*})$, shows the opposite dependence. Numerical simulations suggest that the trends also hold in the case when EEs are of finite volume (Fig. S2). To clarify these trends, we plot $\beta U_{\rm eff}^{\alpha\alpha/\alpha\gamma}(r^{*})$ as functions of $\lambda$ at the strong binding limit according to Eqs. 1 and 12, shown as the black solid and dashed curves in Fig. 4. We note that at this limit, for an isolated PAE $i$, the number of EEs that stay in adsorbed states $M_{i}$ decreases as EE valency increases since fewer EEs are enough to saturate the receptors on the PAE. Moreover, the effective pair interaction in the system mainly depends on $Q$ with $\beta U_{\rm eff}\approx-Q$. For the $\alpha\gamma$ interaction, the bridging of EEs is determined by the adsorption on both particles, where adsorption promotes bridging with $Q_{ij}\sim M_{i}M_{j}$. Therefore, as $\lambda$ increases, both the numbers of EEs in the adsorbed state and bridging state decrease, leading to a weaker effective EE-mediated interaction. Differently, for the $\alpha\alpha$ interaction, we have $Q_{ij}\sim-M_{i}-M_{j}$, indicating that bridging competes with adsorption. Thus, increasing the $\lambda$ decreases the number of EEs in the adsorbed state while increasing the number of EEs in the bridging state, resulting in a stronger effective PAE interaction. At finite binding energies, the effect of increasing $\lambda$ also depends on the types of PAEs involved. As shown in Fig. 4, for $\alpha\alpha$ interactions, increasing the EE valency leads to stronger interactions that eventually converge, similar to the behavior in the strong binding limit. In contrast, the situation is more complex for $\alpha\gamma$ interactions. At relatively weak binding energies (e.g., $\beta\Delta G=-10$), increasing $\lambda$ strengthens the PAE-PAE interaction; at strong binding energies (e.g., $\beta\Delta G=-14$), increasing the EE valency weakens the PAE-PAE interaction; and at intermediate binding energies (e.g., $\beta\Delta G=-12$), the dependence of the strength of the PAE-PAE interaction on $\lambda$ is non-monotonic. The reason for this behavior is that as the binding energy becomes stronger or the valency becomes larger, receptors on the PAEs are more easily saturated. Consequently, the effective interaction $\beta U_{\mathrm{eff}}^{\alpha\gamma}(r^{*})$ more closely approaches that of the strong binding limit.
@@
Furthermore, interestingly, we observe that with increasing $\lambda$, indicated by the gray horizontal line in Fig. 4, the effective PAE interactions unexpectedly converge and become independent of the binding energy and the type of pairs. At this limit, we also have $\bar{n}_{i}\to 0$, similar to the strong binding limit according to Eq. 8. Thus, $\beta F_{\rm att}$ in Eq. 12 holds with the effective binding strength (see Supplementary Material Sec. III.B [32])
@@
This indicates that increasing the valency has less and less impact on effective interactions.
@@
As all interactions converge at large $\lambda$, one can find that there exists an optimal valency $\lambda^{*}$ at which the difference between strengths of effective interactions among different PAEs $|\beta U_{\rm eff}^{\alpha\gamma}(r^{*})-\beta U_{\rm eff}^{\alpha\alpha}(r^{*})|$ maximizes, and $\lambda^{*}$ decreases with decreasing the binding free energy $\Delta G$ (Fig. 4). This optimal $\lambda$ maximizes the disparity between targeted ($\alpha\gamma$) and non-targeted ($\alpha\alpha$) PAE interactions, and our results suggest that at the strong binding limit ($\Delta G\to-\infty$), the optimal valency of EE for mediating specific binding between targeted PAEs is $\lambda^{*}=1$. However, we note that in experiments of designed assembly, the binding strength is normally not very strong, as strong binding would make it hard for the system to reach the designed state in equilibrium. We note that up to 10% polydispersity in the number of grafted strands shows less than $0.3k_{B}T$ deviations in the pair potentials and leaves the optimal valency $\lambda^{*}$ essentially unchanged (Fig. S3), indicating that our design rules are robust to the level of polydispersity reported experimentally.
4 Discussion
@@
To conclude, we have studied the self-assembly of a minimal binary system of PAE-EEs, where EEs act as multivalent linkers with two types of ssDNA ends complementary to those on PAEs. We derived an explicit formula for the effective PAE-PAE interaction mediated by EEs with arbitrary valency $\lambda$, which quantitatively agrees with simulations incorporating explicit EEs. The formulas can be obtained by generalizing the approach in Ref [15, 16] to linkers whose multivalency $\lambda>1$. Here we focus on how the multivalency of linkers influences the effective interaction and phase behaviors in the system. Our analysis reveals an anomalous dependence of the effective PAE-PAE interactions on $\lambda$. Specifically, at the strong binding limit, we found that increasing $\lambda$ strengthens the interaction between PAEs of the same type but weakens that between different types, which converges to the same value in the large valency limit. We identified an optimal valency $\lambda^{*}$ that maximizes the effective interaction difference between the targeted and non-targeted PAE pairs, which provides a practical design principle for the designed self-assembly of PAE-EE systems. The design principle can be implemented directly with current DNA-nanotechnology protocols. Orthogonal 6–8 mer sticky ends that satisfy the required $\Delta G$ can be generated with standard sequence–design software. EE valency is controlled experimentally by varying the density of thiolated oligonucleotides during functionalisation [24]. Moreover, our explicit formula can be directly used in MC simulation to efficiently sample the thermodynamic and structural information of the system. The same saddle-point scheme can be applied to ternary or more complex mixtures by enlarging the binding matrix, and no additional physics is required.
Beyond motivating experimental work towards the design of complex DNA structures, we provide the tools for modeling and analyzing soft matter and biological systems with multivalent linkers, e.g., polyelectrolyte [35, 36], hydrogels [37, 38], vitrimers[39, 40], multivalent drug design [41, 42], and cell adhesion and signaling [43]. While our receptor-mobility model already includes many-body effects among PAEs, higher-order bridges in which a single EE binds three or more particles are not considered here and will be addressed in future work. Moreover, kinetic pathways or possible arrest often accompany short-range attractive colloids [44, 45, 46], and coupling our multivalent-linker framework to real-time kinetic analyses is a promising avenue for future study.
Data availability statement
All data that support the findings of this study are included within the article (and any supplementary files). All scripts used to compute the PAE–EE pair interactions and to run the many-body Monte Carlo simulations are openly available under the MIT licence at https://github.com/xiaxiuyang/multivalent_linker_mediated_interaction.
@@
This work was financially supported by the Academic Research Fund from the Singapore Ministry of Education (RG151/23 and MOE2019-T2-2-010) and the National Research Foundation, Singapore, under its 29th Competitive Research Program (CRP) Call (NRF-CRP29-2022-0002). X. X. acknowledges support from the Alexander von Humboldt-Stiftung.